# Vision Transformers on the Edge: A Comprehensive Survey of Model Compression and Acceleration Strategies

### v1

@@ Abstract
In recent years, vision transformers (ViTs) have emerged as powerful and promising techniques for computer vision tasks such as image classification, object detection, and segmentation. Unlike convolutional neural networks (CNNs), which rely on hierarchical feature extraction, ViTs treat images as sequences of patches and leverage self-attention mechanisms. However, their high computational complexity and memory demands pose significant challenges for deployment on resource-constrained edge devices. To address these limitations, extensive research has focused on model compression techniques and hardware-aware acceleration strategies. Nonetheless, a comprehensive review that systematically categorizes these techniques and their trade-offs in accuracy, efficiency, and hardware adaptability for edge deployment remains lacking. This survey bridges this gap by providing a structured analysis of model compression techniques, software tools for inference on edge, and hardware acceleration strategies for ViTs. We discuss their impact on accuracy, efficiency, and hardware adaptability, highlighting key challenges and emerging research directions to advance ViT deployment on edge platforms, including graphics processing units (GPUs), tensor processing units (TPUs), and field-programmable gate arrays (FPGAs). The goal is to inspire further research with a contemporary guide on optimizing ViTs for efficient deployment on edge devices.
@@ 1 Introduction
Following the tremendous success of transformers in NLP, researchers adapted their architecture for computer vision (CV), leading to the development of vision transformers (ViTs) [4]. Unlike convolutional neural networks (CNNs) [5], which rely on hierarchical feature extraction, ViTs model visual data by processing images as sequences of patch embeddings, enabling global context modeling via self-attention. Since the introduction of ViTs [4], research interest in ViT-based models has grown exponentially, as reflected in the increasing number of publications each year (Figure 1a). This surge in publications highlights ViTs’ dominance in CV tasks, driven by their state-of-the-art (SOTA) performance across various tasks, including image classification [4, 6], object detection [7, 8], and segmentation [9, 10].
While ViT-based models have demonstrated significant capabilities, the substantial size of these models presents major challenges for practical deployment. For instance, ViT-Huge includes over 632M parameters [4] and recently extended to 22B parameters [11], demanding extensive computational resources. These memory and processing requirements make direct deployment on resource-constrained edge devices impractical without optimization. To overcome these limitations, researchers have explored various model compression techniques to reduce computational overhead while preserving performance. As ViTs continue to gain prominence in CV tasks (Figure 1a), there has been a parallel increase in research focused on optimizing their efficiency through compression techniques (Figure 1b). Techniques such as pruning [12, 13], quantization [14, 15], and knowledge distillation (KD) [16, 17] on ViT have gained traction, offering solutions to reduce model size, improve inference speed, and lower power consumption without significantly compromising accuracy.
@@
However, compression techniques alone are often insufficient to meet real-time applications’ stringent latency and throughput requirements on edge devices. To achieve uninterrupted inference in resource-constrained edge devices, acceleration techniques optimize ViT execution by leveraging hardware-aware optimizations, efficient non-linear operations, and efficient resource allocations. These approaches address the inherent on-device computational bottlenecks of ViTs, such as the quadratic complexity of self-attention and the inefficiencies in processing patch embeddings. Recent advancements in accelerating techniques, including the use of specialized or custom accelerators (e.g., graphics processing units (GPUs), tensor processing units (TPUs), and field-programmable gate arrays (FPGAs)) and optimized libraries (e.g., TensorRT), have further expanded the possibilities for accelerating ViTs on edge devices. By bridging the gap between model-level optimizations and hardware-specific execution, software-hardware (SW-HW) co-design also plays a pivotal role in deploying ViTs on devices [19, 20, 21]. Compression techniques, optimized software tools, and hardware-aware acceleration strategies [22, 21, 23] provide a pathway toward efficient, low-latency ViT inference, unlocking new possibilities for autonomous systems, mobile vision applications, and real-time processing on edge devices.
This survey provides a comprehensive review of both model compression and acceleration strategies tailored for ViT, with a particular focus on their applicability to edge devices such as GPUs, central processing units (CPUs), FPGAs, and application-specific integrated circuits (ASICs). We systematically categorize and analyze the latest advancements in pruning, quantization, knowledge distillation, and hardware-aware optimizations. Furthermore, we explore emerging acceleration techniques, which aim to reduce latency and improve energy efficiency. By synthesizing insights from a broad range of studies, this survey serves as a valuable resource for researchers and practitioners seeking to deploy ViTs on edge devices.
@@ 1.1 Motivations and Contribution
ViTs have revolutionized CV tasks, achieving SOTA performance across tasks such as image classification, object detection, and segmentation. However, their high computational cost, memory footprint, and energy consumption present significant challenges for deployment on resource-constrained edge devices. While various optimization techniques exist for ViTs, a comprehensive review that unifies ViT-focused model compression, software tools, evaluation metrics, and hardware acceleration strategies for edge deployment remains underexplored. Existing surveys typically address these aspects in isolation, lacking a holistic analysis that connects them. Table 1 compares existing surveys on ViT model compression and acceleration techniques.
@@
Our Survey ✓✓ ✓✓ ✓✓ ✓✓
This survey addresses this gap by systematically analyzing model compression techniques (pruning, quantization, knowledge distillation) and hardware-aware acceleration strategies (efficient attention mechanisms, SW-HW co-design, FPGA optimizations, etc.). By analyzing insights from a broad range of studies, this survey serves as a valuable resource for researchers and practitioners seeking to deploy ViTs on edge devices. The main contributions of our survey are as follows:
@@
    We systematically categorize and analyze pruning, quantization, and KD to optimize ViTs in resource-constrained environments while maintaining accuracy.
@@
    We investigate current tools for efficient inference and hardware-aware accelerating techniques to enhance ViTs inference efficiency across edge platforms like GPUs, FPGAs, and TPUs.
@@ 1.2 Literature Collection and Organizations
Our literature search was conducted across major academic databases, including Google Scholar, IEEE Xplore, arXiv, and the ACM Digital Library, to ensure comprehensive coverage of relevant research. We utilized targeted search queries with keywords such as vision transformer, acceleration techniques, edge devices, software-hardware co-design, pruning, quantization, and knowledge distillation to identify studies relevant to this survey. A total of 170 papers were collected, with works published up to January 2025 considered for inclusion. Figure 2 illustrates the key concepts discussed in this survey. However, the final selection of papers was determined based on their relevance to key research questions. A paper was included in this survey if it addressed the following criteria:
@@
    Does the paper explore the integration of ViTs with hardware-aware optimizations, including software-hardware co-design strategies?
The remainder of the survey is organized as follows. Section 2 presents an in-depth discussion on model compression techniques, including pruning, quantization, and knowledge distillation, which enhance ViT efficiency while preserving performance across various computer vision tasks. Following this, Section 3 explores an overview of current software tools, optimization frameworks, and evaluation metrics designed for efficient edge inference across different edge devices. Section 4 delves further into hardware-aware accelerating techniques, focusing on optimizations for non-linear operations (e.g., softmax, GELU, and LayerNorm) and current SOTA SW-HW co-design techniques and provides a comprehensive performance comparison of the SOTA techniques. Furthermore, Section 5 discusses key challenges and future research directions, identifying multiple avenues for advancing ViT acceleration and deployment on edge devices. Finally, we conclude this paper in Section 6.
Figure 2: Key concepts discussed in this survey.
@@ 2 Model Compression
Model compression is a key technique for deploying a model on edge devices with limited computational power and memory while maintaining model performance regarding accuracy, precision, and recall. It mainly focuses on lowering latency or reducing memory and energy consumption during inference. However, model compression on ViT still needs extensive exploration due to complex architecture and high resource usage tendencies. In this section, we will discuss prominent compression techniques for ViT models: pruning (Section 2.1), knowledge distillation (Section 2.2), and quantization (Section 2.3).
@@ 2.1 Pruning
Pruning is used for reducing both memory and bandwidth. Most of the initial pruning techniques based on biased weight decay [28], second-order derivatives  [29], and channels  [30]. Early days pruning techniques reduce the number of connections based on the hessian of the loss function [31, 32]. In general, pruning removes redundant parameters that do not significantly contribute to the accuracy of results. The pruned model has fewer edges/connections than the original model. Most early pruning techniques are like brute force pruning, where one needs to manually check which weights do not cause any accuracy loss. Pruning techniques in deep learning became prominent post-2000 as neural networks (NNs) grew in size and complexity. The following subsections will discuss different pruning types and recent pruning techniques applied to ViT-based models.
@@ 2.1.1 Types
Unstructured vs Structured Pruning Unstructured pruning removes individual weights or parameters from the network based on certain criteria. Unstructured pruning can result in highly sparse networks. However, it often does not lead to computational efficiency as the sparsity is not aligned with the memory access patterns or computational primitives of hardware accelerators. However, structured pruning applies to the specific components of the network, especially in layers, neurons, or channels. Structured pruning leads to more hardware-friendly sparsity patterns but often at the cost of higher accuracy loss. Cai et al.  [33] proposed a two-stage coarse-grained/fine-grained structured pruning method based on top-K sparsification and reduces 60% overall computation in the embedded NNs. In a recent survey [34], He et al. discussed a range of SOTA structured pruning techniques, covering topics such as filter ranking methods, regularization methods, dynamic execution, neural architecture search (NAS), the lottery ticket hypothesis, and the applications of pruning.

### v2

@@ Abstract
In recent years, vision transformers (ViTs) have emerged as powerful and promising techniques for computer vision tasks such as image classification, object detection, and segmentation. Unlike convolutional neural networks (CNNs), which rely on hierarchical feature extraction, ViTs treat images as sequences of patches and leverage self-attention mechanisms. However, their high computational complexity and memory demands pose significant challenges for deployment on resource-constrained edge devices. To address these limitations, extensive research has focused on model compression techniques and hardware-aware acceleration strategies. Nonetheless, a comprehensive review that systematically categorizes these techniques and their trade-offs in accuracy, efficiency, and hardware adaptability for edge deployment remains lacking. This survey bridges this gap by providing a structured analysis of model compression techniques, software tools for inference on edge, and hardware acceleration strategies for ViTs. We discuss their impact on accuracy, efficiency, and hardware adaptability, highlighting key challenges and emerging research directions to advance ViT deployment on edge platforms, including graphics processing units (GPUs), application-specific integrated circuit (ASICs), and field-programmable gate arrays (FPGAs). The goal is to inspire further research with a contemporary guide on optimizing ViTs for efficient deployment on edge devices.
@@ 1 Introduction
Following the tremendous success of transformers in NLP, researchers adapted their architecture for computer vision (CV), leading to the development of vision transformers (ViTs) [4]. Unlike convolutional neural networks (CNNs) [5], which rely on hierarchical feature extraction, ViT model processes images as sequences of patch embeddings, enabling global context modeling via self-attention. Since the introduction of ViTs [4], research interest in ViT-based models has grown exponentially, as reflected in the increasing number of publications each year (see Figure 1a). This surge in publications highlights ViTs’ dominance in CV tasks, driven by their state-of-the-art (SOTA) performance across various tasks, including image classification [4, 6], object detection [7, 8], and segmentation [9, 10].
While ViT-based models have demonstrated significant capabilities, the substantial size of these models presents major challenges for practical deployment. For instance, ViT-Huge includes over 632M parameters [4] and has recently been extended to 22B parameters [11], demanding extensive computational resources. These memory and processing requirements make direct deployment on resource-constrained edge devices impractical without optimization. To overcome these limitations, researchers have explored various model compression techniques to reduce computational overhead while preserving performance. As ViTs continue to gain prominence in CV tasks (Figure 1a), there has been a parallel increase in research focused on optimizing their efficiency through compression techniques (Figure 1b). Techniques such as pruning [12, 13], quantization [14, 15], and knowledge distillation (KD) [16, 17] on ViT have gained traction, offering solutions to reduce model size, improve inference speed, and lower power consumption without significantly compromising accuracy.
@@
However, compression techniques alone are often insufficient to meet real-time applications’ stringent latency and throughput requirements on edge devices.
Although model compression techniques reduce computational overhead, they are often insufficient to ensure real-time performance on edge devices due to hardware-specific constraints. Effective deployment of ViT requires tailored acceleration strategies that align with the architectural characteristics of each hardware platform. These strategies optimize execution by addressing ViTs’ core computational bottlenecks—such as self-attention’s quadratic complexity and patch embedding inefficiencies—through hardware-aware optimizations, efficient non-linear operations, and resource-efficient computation. These approaches address the inherent on-device computational bottlenecks of ViTs, such as the quadratic complexity of self-attention and the inefficiencies in processing patch embeddings. Recent advancements in accelerating techniques, including the use of specialized or custom accelerators and optimized libraries (e.g., TensorRT), have further expanded the possibilities for accelerating ViTs on edge devices (e.g., graphics processing units (GPUs), application-specific integrated circuits (ASICs), and field-programmable gate arrays (FPGAs)). By bridging the gap between model-level optimizations and hardware-specific execution, software-hardware (SW-HW) co-design also plays a pivotal role in deploying ViTs on edge devices [19, 20, 21]. Compression techniques, optimized software tools, and hardware-aware acceleration strategies [22, 21, 23] provide a pathway toward efficient, low-latency ViT inference, unlocking new possibilities for different application domains such as autonomous driving, mobile vision applications, and medical imaging on edge devices.
This survey provides a comprehensive review of both model compression and acceleration strategies tailored for ViT, focusing on their applicability to edge devices such as GPUs, central processing units (CPUs), FPGAs, and ASICs. We systematically categorize and analyze the latest advancements in pruning, quantization, KD, and hardware-aware optimizations. Furthermore, we explore emerging acceleration techniques, which aim to reduce latency and improve energy efficiency. By synthesizing insights from a broad range of studies, this survey serves as a valuable resource for researchers and practitioners seeking to deploy ViTs on edge devices.
@@ 1.1 Motivations and Contribution
ViTs have revolutionized CV tasks, achieving SOTA performance across tasks such as image classification, object detection, and segmentation. However, their high computational cost, memory footprint, and energy consumption present significant challenges for deployment on resource-constrained edge devices. While various optimization techniques exist for ViTs, a systematic survey that collectively examines ViT-focused model compression, software tools, evaluation metrics, and hardware acceleration strategies for edge deployment remains limited. Existing surveys often treat these aspects independently, lacking a single resource that presents them together in a structured manner tailored to edge deployment. Table 1 compares existing surveys on ViT model compression and acceleration techniques.
@@
Our Survey 2025 ✓✓ ✓✓ ✓✓ ✓✓
This survey addresses this gap by systematically analyzing model compression techniques (pruning, quantization, KD) and hardware-aware acceleration strategies (efficient attention mechanisms, SW-HW co-design, FPGA optimizations, etc.). By analyzing insights from a broad range of studies, this survey serves as a valuable resource for researchers seeking to deploy ViTs on edge devices. The main contributions of our survey are as follows:
@@
    We systematically categorize and analyze model compression techniques such as pruning, quantization, and KD to optimize ViTs for deploying on edge devices.
@@
    We investigate current software tools for efficient inference and hardware-aware accelerating techniques to enhance inference efficiency across edge platforms like GPUs, FPGAs, and ASICs for ViT models.
@@ 1.2 Literature Collection and Organizations
Our literature search was conducted across major academic databases, including Google Scholar, IEEE Xplore, arXiv, and the ACM Digital Library, to ensure comprehensive coverage of relevant research. We utilized targeted search queries with keywords such as vision transformer, acceleration techniques, edge devices, software-hardware co-design, pruning, quantization, and knowledge distillation to identify studies relevant to this survey. A total of 170 papers were collected, with works published up to January 2025 considered for inclusion. Figure 2 presents a structured overview of the selected literature, categorizing each referenced work according to the core research themes and enabling tools addressed in this survey. Papers were included based on how well they aligned with our key research questions. A paper was selected if it met one or more of the following criteria:
@@
    Does the paper explore the integration of ViTs with hardware-aware optimizations, including SW-HW co-design strategies?
The remainder of the survey is organized as follows. Section 2 presents an in-depth discussion on model compression techniques, including pruning, quantization, and KD, which enhance ViT efficiency while preserving performance across various CV tasks. Following this, Section 3 explores an overview of current software tools, optimization frameworks, and evaluation metrics designed for efficient edge inference across different edge devices. Section 4 further explores hardware-aware acceleration techniques, focusing on optimizations for non-linear operations (e.g., softmax, GELU, and LayerNorm) and recent SOTA SW-HW co-design strategies, along with a comparative analysis of their performance. Furthermore, Section 5 discusses key challenges and future research directions, identifying multiple avenues for advancing ViT acceleration and deployment on edge devices. Finally, we conclude this paper in Section 6.
Figure 2: Core components and tools are covered in this survey. This survey focuses on three main parts—model compression techniques (pruning, KD, and quantization), accelerating techniques (using FPGA, GPU, and ASIC), and associated software tools (toolkits, libraries, and inference engines)—to provide a comprehensive understanding of efficient ViT deployment.
@@ 2 Model Compression
Model compression is a key technique for deploying a model on edge devices with limited computational power and memory while maintaining accuracy. It enables lowering inference latency as well as reducing memory and energy consumption. However, model compression on ViT still needs extensive exploration due to its complex architecture and high resource usage tendencies. In this section, we will discuss prominent compression techniques for ViT models: pruning (Section 2.1), KD (Section 2.2), and quantization (Section 2.3).
@@ 2.1 Pruning
Pruning is used to reduce both memory and bandwidth. Most of the initial pruning techniques based on biased weight decay [28], second-order derivatives [29], and channels [30]. Early days pruning techniques reduce the number of connections based on the hessian of the loss function [31, 32]. In general, pruning removes redundant parameters that do not significantly contribute to the accuracy of results. The pruned model has fewer edges/connections than the original model. Most early pruning techniques are like brute force pruning, where one needs to manually check which weights do not cause any accuracy loss. Pruning techniques in deep learning became prominent post-2000 as neural networks (NNs) grew in size and complexity. The following subsections will discuss different pruning types and recent techniques applied to ViT-based models.
@@ 2.1.1 Types
Unstructured vs Structured Pruning Unstructured pruning removes individual weights or parameters from the network based on certain criteria. Unstructured pruning can result in highly sparse networks. However, it often does not lead to computational efficiency as the sparsity is not aligned with the memory access patterns or computational primitives of hardware accelerators. However, structured pruning applies to the specific components of the network, especially in layers, neurons, or channels. Structured pruning leads to more hardware-friendly sparsity patterns but often at the cost of higher accuracy loss. Cai et al. [33] proposed a two-stage coarse-grained/fine-grained structured pruning method based on top-K sparsification, reducing 60% overall computation in the embedded NNs. In a recent survey [34], He et al. discussed a range of SOTA structured pruning techniques, covering topics such as filter ranking methods, regularization methods, dynamic execution, neural architecture search (NAS), the lottery ticket hypothesis, and the applications of pruning.
Figure 3: Static Pruning vs Dynamic Pruning
In contrast, dynamic pruning applies during runtime based on specific criteria, such as the importance of neurons or weights. One of the significant drawbacks of static pruning is that it relies on a fixed pruning schedule and rate, which is determined before training begins. That means that static pruning does not adjust to the network’s learning progress or the changing importance of neurons or weights during training. Dynamic pruning evaluates and adjusts the pruning criteria during training based on real-time importance assessments to overcome the limitation of static pruning. Most of the recent pruning techniques use dynamic pruning techniques [36, 13, 37] to get the accuracy without loss of any information.
Table 2: Categorization of ViT pruning methods across key perspectives. Many methods span multiple categories.
Method Importance-Based Token Pruning Structured Cascade
VTP [38]
Patch Slimming [39] Top-down
WDPruning [40]
SP-ViT [37]
UVC [41]
CP-ViT [36]
NViT [42]
S2ViT [43]
VTC-LFC [44]
2.1.2 Pruning Techniques for Vision Transformer
Pruning methods for ViT-based models remain an underexplored area, with only a handful of studies in recent years. Current ViT pruning methods can be mainly divided into importance-based, token pruning, structure pruning, cascading pruning (combining multiple pruning techniques for each module), and miscellaneous methods. Note that many of these pruning approaches overlap across categories. For instance, patch slimming [39] and CP-ViT [36] apply both importance-based and structured token pruning, while CP-ViT [36] and VTC-LFC [44] also follow a cascade strategy. Table 2 summarizes the categorization across different pruning approaches. This section provides a brief overview of the current SOTA pruning techniques for ViTs.
Structure Pruning Recently, many studies on ViT pruning have embraced structure pruning techniques to optimize model efficiency. As discussed in the previous subsections, importance-based pruning and token pruning can also be considered as structured pruning. Additionally, Yu et al. [41] proposed a structure pruning in a ViT named UVC, where they pruned the number of heads and dimensions inside each layer. Experiments of this paper were conducted in various ViT models (e.g., DeiT-Tiny and T2T-ViT-14) on ImageNet-1k [45] datasets. DeiT-Tiny [46] cut down to 50% of the original floating-point operations per second (FLOPS) while not dropping accuracy much in this study. Another study [47] proposed structure pruning on multi-head self-attention (MSA) [1] and feedforward neural network (FFN) [1] by removing unnecessary parameter groups. Other studies on structure pruning named NViT [42] proposed hessian-based iterative structure pruning criteria comparable across all layers and structures. Moreover, it incorporated latency-aware regularization techniques to reduce latency directly. Another study on the structure pruning in ViT called S2ViT [43] removed submodules like self-attention heads by manipulating the weight, activations & gradients. Additionally, current structure pruning methods on ViTs utilize either importance-based or learnable token selector pruning.
Figure 4: The overview of importance-based pruning techniques for ViTs. VTP [38] prunes dimensionality in the multi-head self-attention (MSA) and multi-layer perception (MLP) modules using important scores. WDPruning [40] applies a binary mask ($N$) to the MSA module (a), followed by pruning of (b) attention heads and (c) linear projection channels. Patch Slimming [39] removes redundant patches in a top-down, layer-wise manner.
Importance-based Pruning One of the significant challenges in importance-based pruning for ViTs lies in determining effective thresholds to identify which components—such as channels, attention heads, blocks, or input patches—should be pruned. Different pruning strategies address this by designing tailored scoring criteria and thresholding mechanisms, often customized based on the targeted layer type or pruning granularity. Since a significant portion of ViT’s computational burden stems from the MSA and multi-layer perceptron (MLP) modules [1], targeted pruning of these modules can yield substantial computational savings. Figure 4 summarizes the current pruning techniques on ViT based on different important score measurements. Zhu et al. [38] introduced vision transformer pruning (VTP), one of the first dedicated pruning methods for ViTs that removes dimensions with lower importance scores, achieving a high pruning ratio without sacrificing accuracy. VTP operates in three key steps: 1) L1 sparse regularization is applied during training to identify less significant channels, 2) channel pruning eliminates redundant computations based on thresholds derived from a predefined pruning rate, and (3) finetuning. However, VTP ignores network depth, which is an important dimension for pruning the model. To address this issue, another study [40] (see Figure 4(a–c)) explored width and depth pruning on the MSA module in the transformer models. Block a in Figure 4 illustrates the application of binary masking $N$ within the MSA module. Additionally, Blocks b and c in Figure 4 depict width pruning applied to the linear projection matrices and attention head pruning, respectively. Moreover, leveraging the learnable saliency score approach enables non-uniform sparsity across layers and facilitates progressive block removal, allowing the model to be pruned sequentially during a single training epoch. Furthermore, Tang et al. [39] identified that many input patches exhibit redundancy across transformer layers. Based on this insight, they proposed a layer-by-layer patch-slimming approach that reduces redundant patches in a top-down manner, offering a more adaptive alternative to conventional structured pruning techniques. The authors calculated their importance scores for the final classification feature to identify unimportant patches. The proposed method started by identifying important patches in the last transformer layer and recursively utilized them back to patches in earlier layers.
Figure 5: The overview of token pruning based on SP-ViT [37].
Token Pruning Kong et al. [37] introduced a latency-aware soft token pruning framework, SP-ViT. Figure 5 illustrates the workflow of SP-ViT token pruning on ViT models. The framework was implemented on vanilla transformers such as data-efficient image transformers (DeiT) [46] and swin transformers [6]. The authors proposed a dynamic attention-based multi-head token selector for adaptive instance-wise token selection. Subsequently, they incorporated a soft pruning method that consolidated less informative tokens into a package token instead of entirely discarding them, identified by the selector module. The authors evaluated their proposed method on ImageNet-1k [45] dataset using Swin-S, Swin-S/T [6], PiT-Xs/S [48] models. SP-ViT was subsequently adopted by FPGA-based accelerating frameworks such as HeatViT [20]. While SP-ViT [37] was initially developed and evaluated for image classification, recent studies have extended token pruning techniques to more complex tasks such as object detection [49] and semantic segmentation [50] using lightweight, background-aware ViTs.
Table 3: Top-1 (%) accuracy comparison of pruning techniques across different ViT models. $\downarrow$’ denotes accuracy loss from the baseline ViT model, and $\uparrow$’ denotes the accuracy gain after applying the pruning technique. All results are evaluated on the ImageNet-1k [45] dataset.
Algorithm Method Top-1 Accuracy (%) on ViT Models
DeiT-T DeiT-S DeiT-B Swin-S T2T-ViT-14
Baseline - 72.2 79.8 81.8 83.2 81.5
VTP [38] Dimension-based - -
80.7
($\downarrow$ 1.34%)
WDPruning [40] Width-depth pruning
70.34
($\downarrow$ 2.58%)
78.3
($\downarrow$ 1.88%)
80.8
($\downarrow$ 1.22%)
81.8
($\downarrow$ 1.68%)
MD [51]
Dependency based pruning
by Gaussian-search
79.9
($\uparrow$ 0.13%)
82.3
($\downarrow$ 0.61%)
81.7
($\uparrow$ 0.24%)
SPViT [37]
Pruning based on
token score
82.7
($\downarrow$ 0.60%)
PS-ViT [39]
Layer-wise top
down pruning
72.0
($\downarrow$ 0.28%)
79.4
($\downarrow$ 0.50%)
81.5
($\downarrow$ 0.30%)
81.1
($\downarrow$ 0.49%)
UVC [41] Head/dimension pruning
70.6
($\downarrow$ 2.08%)
78.8
($\downarrow$ 1.25%)
80.6
($\downarrow$ 1.47%)
78.9
($\downarrow$ 3.19%)
NViT [42]
Latency-aware, Hessian-
based pruning
76.2
($\uparrow$ 5.55%)
82.2
($\uparrow$ 3.01%)
83.3
($\uparrow$ 1.83%)
83.0
($\downarrow$ 0.24%)
SAViT [47]
Structural pruning on
MSA attention & FFN
70.7
($\downarrow$ 2.08%)
80.1
($\uparrow$ 0.37%)
81.7
($\uparrow$ 0.12%)
S2ViTE [43]
Removing sub-modules
by manipulating weight,
activation,& gradient
70.1
($\downarrow$ 2.91%)
79.2
($\downarrow$ 0.75%)
82.2
($\uparrow$ 0.49%)
VTC-LFC [44]
Token & channel pruning
using a hyperparameter
71.0
($\downarrow$ 1.66%)
79.6
($\downarrow$ 0.25%)
81.6
($\downarrow$ 0.24%)
CP-ViT [36]
Prune PH-regions in
MSA & FFN
dynamically
71.2
($\downarrow$ 1.39%)
79.0
($\downarrow$ 1.00%)
81.1
($\downarrow$ 0.86%)
  • 1.
    PH-region: Informative Patches and attention Heads identified during cascade pruning.
Cascade Pruning Cascade pruning combines multiple pruning techniques to reduce parameters and giga FLOPS (GFLOPS) while preserving accuracy. A standout method named VTC-LFC [44] aimed to improve the identification of informative channels and tokens in a model, leading to better accuracy preservation. This approach introduced a bottom-up cascade (BCP) pruning strategy that gradually prunes tokens and channels, starting from the first block and advancing to the last. The pruning process is controlled by a hyperparameter called a global allowable drop, ensuring the performance drop remains within an acceptable range. Additionally, BCP guarantees efficient compression without sacrificing model performance by pruning each block and immediately stopping the compression process when the performance drop reaches a predefined threshold. Another cascade ViT pruning  [36] utilized the sparsity for pruning PH-regions (containing patches and heads) in the MSA & FFN progressively and dynamically. The authors conducted experiments on three different types of datasets: ImageNet-1k [45], CIFAR-10 [52], and CIFAR-100 [52].